A load W = 6000N is supported by two ropes one of which makes an angle of

     50o with the horizontal while the other rope makes an angle of 60o with the

     Vertical. Calculate the force exerted by each rope so as to provide

      equilibrium. CALCULATE the Resultant, and Resultant Angle.


1
Expert's answer
2021-05-12T14:58:43-0400

Draw the diagram:



From the diagram, we see that for the weight to remain in equilibrium, two conditions are important: the sum of horizontal components of all forces must zero, the sum of vertical components of all forces must be equal:


"-T_1\cos50°+T_2\cos40°=0,\\
-W+T_1\sin50°+T_2\sin40°=0.\\\space\\
T_1=T_2\frac{\cos40°}{\cos50°},\\\space\\
-W+T_2\frac{\cos40°}{\cos50°}\sin50°+T_2\sin40°=0,\\\space\\
T_2=\frac{W}{\cos40°·\tan50°+\sin40°}=3857\text{ N},\\\space\\
T_1=\frac{W\cos40°}{\cos50°(\cos40°·\tan50°+\sin40°)}=4596\text{ N}."

The angle between the ropes is


"\theta=180°-50°-40°=90°."